# MambaVision: A Hybrid Mamba-Transformer Vision Backbone

### v1

@@ Abstract
We propose a novel hybrid Mamba-Transformer backbone, denoted as MambaVision, which is specifically tailored for vision applications. Our core contribution includes redesigning the Mamba formulation to enhance its capability for efficient modeling of visual features. In addition, we conduct a comprehensive ablation study on the feasibility of integrating Vision Transformers (ViT) with Mamba. Our results demonstrate that equipping the Mamba architecture with several self-attention blocks at the final layers greatly improves the modeling capacity to capture long-range spatial dependencies. Based on our findings, we introduce a family of MambaVision models with a hierarchical architecture to meet various design criteria. For Image classification on ImageNet-1K dataset, MambaVision model variants achieve a new State-of-the-Art (SOTA) performance in terms of Top-1 accuracy and image throughput. In downstream tasks such as object detection, instance segmentation and semantic segmentation on MS COCO and ADE20K datasets, MambaVision outperforms comparably-sized backbones and demonstrates more favorable performance. Code: https://github.com/NVlabs/MambaVision.
@@ 1 Introduction
[…]
Figure 1: Top-1 accuracy vs. image throughput on ImageNet-1K. All measurements conducted on A100 GPU with batch size of 128. MambaVision achieves a new SOTA Pareto front.
During the recent years, Transformers [1] have become the de facto architecture in different domains including in computer vision, natural language processing, speech processing and robotics. In addition, the versatility of the Transformer architecture, primarily due to its attention mechanism, and its flexibility makes it highly suitable for multimodal learning tasks, where integrating and processing information from various modalities is essential. Despite these benefits, the quadratic complexity of the attention mechanism with respect to sequence length make Transformers computationally expensive to train and deploy. Recently, Mamba [2] proposed a new State Space Model (SSM) that achieves linear time complexity and outperforms or matches Transformers in different language modeling tasks [2]. The core contribution of Mamba is a novel selection mechanism that enables efficient input-dependent processing of long sequences with hardware-aware considerations.
@@
Recently, a number of Mamba-based backbones [3, 4] have also been proposed to leverage the strengths of its SSM formulation in vision tasks such as image classification and semantic segmentation. However, the Mamba’s autoregressive formulation, while effective for tasks requiring sequential data processing, faces limitations in computer vision tasks that benefit from a full receptive field: (1) Unlike sequences where order matters, image pixels do not have a sequential dependency in the same way. Instead, spatial relationships are often local and need to be considered in a more parallel and integrated manner. Hence, this results in inefficiency for processing spatial data (2) an autoregressive model like Mamba processes data step-by-step, limiting its ability to capture and utilize global context in one forward pass. In contrast, vision tasks often require understanding the global context to make accurate predictions about local regions.
Vision Mamba (Vim) [3] and others have proposed modifications such as bidirectional SSMs to address lack of global context and spatial understanding. While bidirectional SSMs have the potential to capture more comprehensive context, they introduce significant latency due to the need to process the entire sequence before making predictions. Additionally, the increased complexity can lead to challenges in training, risk of overfitting, and may not always result in better accuracy. Due to these pitfalls, backbones with Vision Transformer (ViT) and Convolutional Neural Network (CNN) architectures still outperform best Mamba-based vision models on different vision tasks.
@@
In this work, we systematically re-design the Mamba block to make it more suitable for vision tasks. We propose a hybrid architecture that consists of our proposed formulation (i.e. MambaVision Mixer and MLP) as well as Transformer blocks. Specifically, we study different integration patterns such as adding the Transformer blocks in an iso-parameter manner to earlier, middle and final layers as well as every $l$ layers. Our analysis shows that leveraging several self-attention blocks at the final stages can significantly enhance the capability to capture global context and long-range spatial dependencies. As shown in Sec. 5, using a hybrid architecture also results in a higher image throughput in comparison to both pure Mamba or ViT-based models.
We introduce the MambaVision model which consists of a multi-resolution architecture and leverages CNN-based residual blocks for fast feature extraction of larger resolution features. As shown in Fig. 1, the MambaVision achieves a new SOTA Pareto front in terms of ImageNet-1K Top-1 accuracy and image throughput, outperforming Mamba, CNN and ViT-based models, sometimes by a significant margin. In downstream tasks such as Object detection and instance segmentation as well as semantic segmentation, models with MambaVision backbones outperform comparably-sized counterparts on MS COCO and ADE20 datasets, respectively. Hence, it validates the effectiveness and versatility of MambaVision as an efficient backbone.
@@
  • We present a systematic investigation of integration patterns for Mamba and Transformer blocks and demonstrate that incorporating self-attention blocks at the final stages significantly improves the model’s ability to capture global context and long-range spatial dependencies.
  • We introduce MambaVision which is a novel hybrid Mamba Transformer model. The hierarchical MambaVision achieves a new SOTA Pareto front on ImageNet-1K dataset in terms of Top-1 and image throughput tradeoff.
@@ 2 Related work
ViT. The Vision Transformer (ViT) [5] emerged as a promising alternative to CNNs, leveraging self-attention layers to provide an enlarged receptive field. However, ViTs initially lacked some of the inherent advantages of CNNs, such as inductive biases and translation invariance, and they required large-scale training datasets to achieve competitive performance. To address these limitations, the Data-efficient Image Transformer (DeiT) [6] introduced a distillation-based training strategy that significantly improved classification accuracy, even with smaller datasets. Building on this, the LeViT [7] model proposed a hybrid approach, incorporating redesigned MLP and self-attention modules optimized for fast inference, enhancing both efficiency and performance. Furthermore, the Cross-covariance Image Transformer (XCiT) [8] introduced a transposed self-attention mechanism, effectively modeling interactions between feature channels and improving the model’s ability to capture complex patterns in the data. The Pyramid Vision Transformer (PVT) [9] adopted a hierarchical structure with patch embedding at the start of each stage and spatial dimension reduction, thereby improving computational efficiency. Similarly, Swin Transformers [10] proposed a layered architecture where self-attention is computed within local windows, which are shifted to enable interaction between regions, balancing local and global contexts. The Twins Transformer [11] featured spatially separable self-attention that significantly enhanced efficiency. In addition, the Focal Transformer [12] utilized focal self-attention for capturing fine-grained details of long-range spatial interactions.
Mamba. Since the introduction of Mamba, a number of efforts have been proposed to leverage its capability for vision applications. Specifically, Vim [3] proposed to use a bidirectional SSM formulation, with the same Mamba formulation, in which tokens were processed in both forward and backward directions in order to capture more global context and improve spatial understanding. However, bidirectional encoding increases the computational load which can slow down training and inference times. In addition, combining information from multiple directions effectively to form a coherent global understanding is challenging as some global context may be lost in the process. As opposed to Vim, our proposed MambaVision uses a single forward pass with a redesigned Mamba block that can capture both short and long-range information and outperform it significantly both in terms of ImageNet Top-1 accuracy and throughput.
EfficientVMamba [4] proposed an atrous-based selective scan with a skip sampling approach to efficiently extract global spatial dependencies. EfficientVMamba also uses a hierarchical architecture, composed of SSM and CNN-based blocks, in which SSMs are used for larger input resolutions to better capture global context while CNNs are used for lower resolutions. In contrast to EfficientVMamba, MambaVision uses CNNs in higher resolutions for faster feature extraction while using both SSMs and self-attention in lower resolution to capture fine-grained details in both short and long-range spatial dependencies. Our proposed MambaVision also outperforms EfficientVMamba by a significant margin in terms of both Top-1 accuracy and image throughput.
In addition, VMamba [13] introduced a generic Mamba-based vision backbone with a Cross-Scan Module (CSM) which facilitates 1D selective scan with an enlarged global receptive field. Specifically, the CSM modules adopt a four-way selective scan methodology (i.e. upper-left and lower-right to opposite directions) to integrate information from all surrounding tokens and capture more global context. In addition, VMamba makes architectural changes such as using depth-wise convolutions and a hierarchical multi-resolution structure. Although the design of the CSM module is more suited for vision tasks, the receptive field is still limited by the cross-scan paths. In comparison to VMamba, the design of our proposed MambaVision mixer is simpler and can capture both short and long-range dependencies. MambaVision also uses CNN-based layer for fast feature extraction as opposed to VMamba which uses the same block structures in all stages. Furthermore, the MambaVision models outperform VMamba counterparts while having significantly higher image throughput.
3 Methdology
@@ 3.1 Macro Architecture
In this section, we introduce MambaVision which is our proposed novel architecture with SOTA performance on ImageNet-1K dataset. As illustrated in Fig. 2, MambaVision has a hierarchical architecture consisting of $4$ different stages. The first two stages consist of CNN-based layers for fast feature extraction at higher input resolutions, while stage $3$ and $4$ include the proposed MambaVision and Transformer blocks. Specifically, given an image of size $H\times W\times 3$, the input is first converted into overlapping patches with size $\frac{H}{4}\times\frac{W}{4}\times C$ and projected into a $C$ dimensional embedding space by the stem which consists of two consecutive $3\times 3$ CNN layers with stride of $2$. The downsampler in between stages consists of a batch normalized $3\times 3$ CNN layer with stride $2$ which reduces the image resolution by half. Furthermore, the CNN blocks in stages $1$ and $2$ follow a generic residual block formulation according to the following
Figure 2: The architecture of hierarchical MambaVision models. The first two stages use residual convolutional blocks for fast feature extraction. Stage 3 and 4 employ both MambaVision and Transformer blocks. Specifically, given $N$ layers, we use $\frac{N}{2}$ MambaVision and MLP blocks which are followed by additional $\frac{N}{2}$ Transformer and MLP blocks. The Transformer blocks in final layers allow for recovering lost global context and capture long-range spatial dependencies.
@@
GELU and BN denote Gaussian Error Linear Unit activation function [14] and batch normalization [15], respectively.
3.2 Micro Architecture
3.2.1 Mamba Preliminaries
Mamba is an extension of Structured state space sequence models (S4) which is capable of transforming a 1D continuous input $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ via a learnable hidden state $h(t)\in\mathbb{R}^{M}$ with parameters $\bm{A}\in\mathbb{R}^{M\times M}$, $\bm{B}\in\mathbb{R}^{1\times M}$ and $\bm{C}\in\mathbb{R}^{1\times M}$ according to
Discretization
The continuous parameters $\bm{A}$, $\bm{B}$ and $\bm{C}$ in the above formulation are further converted into discrete parameters for better computational efficiency [16]. Specifically, assuming a timescale $\Delta$, a zero-order hold rule can be applied to obtain discrete parameters $\bm{\bar{A}}\in\mathbb{R}^{M\times M}$, $\bm{\bar{B}}\in\mathbb{R}^{1\times M}$ and $\bm{\bar{C}}\in\mathbb{R}^{1\times M}$ according to
$\bar{\bm{A}}=\exp{(\Delta\bm{A})},$ (3)
$\bar{\bm{C}}=\bm{C},$
The Eq. 2 can then be expressed with discrete parameters as
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$ (4)
$y(t)=\bar{\bm{C}}h(t),$
$\begin{split}\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{T-1}\bm{\overline{B}})\\ \quad\bm{y}&=\bm{x}*\bm{\overline{K}},\\ \end{split}$ (5)
Selectivity
3.2.2 Layer Architecture
$\begin{array}[]{l}\hat{{X}}^{n}=\text{Mixer}(\text{Norm}({X}^{n-1}))+{X}^{n-1}\\ {X}^{n}=\text{MLP}(\text{Norm}(\hat{{X}}^{n}))+\hat{{X}}^{n},\\ \end{array}$ (6)
@@ 3.2.2 Layer Architecture
dt = torch.exp(torch.rand(self.dim//2) * (math.log(dt_max) math.log(dt_min)) + math.log(dt_min)
@@
As shown in Fig. 3, we redesigned the original Mamba mixer to make it more suitable for vision tasks. First, we propose to replace the causal convolution with regular convolution, since it limits the influence to one direction, which is unnecessary and restrictive for vision tasks. In addition, we added a symmetric branch without SSM, consisting of an additional convolution and SiLU activation, to compensate for any content lost due to the sequential constraints of SSMs. We then concatenate the output of both branches and project it via a final linear layer. This combination ensures that the final feature representation incorporates both the sequential and spatial information, leveraging the strengths of both branches. We note that the output of each branch is projected into an embedding space with size $\frac{C}{2}$ (i.e. half the size of original embedding dimension) to maintain similar number of parameters to the original block design. Given an input $X_{in}$, the output of MambaVision mixer $X_{out}$is computed according to
$\begin{array}[]{l}{X_{1}}=\text{Scan}(\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}}))))\\ {X_{2}}=\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}})))\\ {X_{out}}=\text{Linear}(\frac{C}{2},C)(\text{Concat}(X_{1},X_{2})),\end{array}$ (7)
$\text{Linear}(C_{in},C_{out})(\cdot)$ denotes a linear layer with $C_{in}$ and $C_{out}$ as input and output embedding dimensions, Scan is the selective scan operation as in [2] and $\sigma$ is the activation function for which Sigmoid Linear Unit (SiLU) [17] is used. In addition, Conv and Concat represent 1D convolution and concatenation operations. In Algorithm 1, we present a PyTorch-like pseudo-code for MambaVision mixer. In general, our proposed modification leads to richer feature representations, better generalization, and improved performance on computer vision tasks. We have also experimentally validated the effectiveness of each of our design choices in Sec.
@@
$Q,K,V$ denote query, key and value respectively and $d_{h}$ is the number of attention heads. Without loss of generality, the attention formulation which can be computed in a windowed manner similar to previous efforts [10, 18]
Table 1: Classification benchmarks on ImageNet-1K [19] dataset. Image throughput is computed on A100 GPU with a batch size of 128.
Model Image Size #Param FLOPs Throughput Top-1

### v2

@@ Abstract
We propose a novel hybrid Mamba-Transformer backbone, MambaVision, specifically tailored for vision applications. Our core contribution includes redesigning the Mamba formulation to enhance its capability for efficient modeling of visual features. Through a comprehensive ablation study, we demonstrate the feasibility of integrating Vision Transformers (ViT) with Mamba. Our results show that equipping the Mamba architecture with self-attention blocks in the final layers greatly improves its capacity to capture long-range spatial dependencies. Based on these findings, we introduce a family of MambaVision models with a hierarchical architecture to meet various design criteria. For classification on the ImageNet-1K dataset, MambaVision variants achieve state-of-the-art (SOTA) performance in terms of both Top-1 accuracy and throughput. In downstream tasks such as object detection, instance segmentation, and semantic segmentation on MS COCO and ADE20K datasets, MambaVision outperforms comparably sized backbones while demonstrating favorable performance. Code: https://github.com/NVlabs/MambaVision
[…]
Figure 1: Top-1 accuracy vs. image throughput comparisons on ImageNet-1K dataset. The MambaVision models achieve a new Pareto front for Top-1 accuracy and image throughput tradeoff. Specifically, MambaVision variants outperform Mamba-based models such as VMamba and Vim, sometimes by a significant margin. For all models, image throughput is measured on an A100 NVIDIA GPU with a batch size of 128.
@@ 1 Introduction
In recent years, Transformers [31] have become the de facto architecture across different domains including computer vision, natural language processing, speech processing, and robotics. The versatility of the Transformer architecture, primarily due to its attention mechanism and flexibility, makes it highly suitable for multimodal learning tasks, where integrating and processing information from various modalities is essential. Despite these benefits, the quadratic complexity of the attention mechanism with respect to sequence length makes Transformers computationally expensive to train and deploy. Recently, Mamba [7] proposed a new State Space Model (SSM) that achieves linear time complexity and outperforms or matches Transformers in different language modeling tasks [7]. The core contribution of Mamba is a novel selection mechanism that enables efficient input-dependent processing of long sequences with hardware-aware considerations. Recently, a number of Mamba-based backbones [40, 25] have been proposed to leverage the strengths of its SSM formulation in vision tasks such as image classification and semantic segmentation. However, the Mamba’s autoregressive formulation, while effective for tasks requiring sequential data processing, faces limitations in computer vision tasks that benefit from a full receptive field: (1) Unlike sequences where order matters, image pixels do not have a sequential dependency in the same way. Instead, spatial relationships are often local and need to be considered in a more parallel and integrated manner. Hence, this results in inefficiency for processing spatial data (2) An autoregressive model like Mamba processes data step-by-step, limiting its ability to capture and utilize global context in one forward pass. In contrast, vision tasks often require understanding the global context to make accurate predictions about local regions. Vision Mamba (Vim) [40] and others have proposed modifications such as bidirectional SSMs to address lack of global context and spatial understanding. While bidirectional SSMs have the potential to capture more comprehensive context, they introduce significant latency due to the need to process the entire sequence before making predictions. Additionally, the increased complexity can lead to challenges in training, risk of overfitting, and may not always result in better accuracy. Due to these pitfalls, backbones with Vision Transformer (ViT) and Convolutional Neural Network (CNN) architectures still outperform best Mamba-based vision models on different vision tasks.
@@
In this work, we systematically re-design the Mamba block to make it more suitable for vision tasks. We propose a hybrid architecture that consists of our proposed formulation (i.e. MambaVision Mixer and MLP) as well as Transformer blocks. Specifically, we study different integration patterns such as adding the Transformer blocks in an iso-parameter manner to earlier, middle, and final layers as well as every $l$ layers. Our analysis shows that leveraging several self-attention blocks at the final stages can significantly enhance the capability to capture global context and long-range spatial dependencies. As shown in Sec. 5, using a hybrid architecture also results in higher image throughput compared to both pure Mamba and ViT-based models. We introduce the MambaVision model which consists of a multi-resolution architecture and leverages CNN-based residual blocks for fast feature extraction of larger resolution features. As shown in Fig. 1, the MambaVision achieves a new SOTA Pareto front in terms of ImageNet-1K Top-1 accuracy and image throughput, outperforming Mamba, CNN, and ViT-based models, sometimes by a significant margin. In downstream tasks such as object detection, instance segmentation, and semantic segmentation, models with MambaVision backbones outperform comparably-sized counterparts on MS COCO and ADE20 datasets, respectively. Hence, it validates the effectiveness and versatility of MambaVision as an efficient backbone.
@@
  • We present a systematic investigation of integration patterns for Mamba and Transformer blocks, and demonstrate that incorporating self-attention blocks at the final stages significantly improves the model’s ability to capture global context and long-range spatial dependencies.
  • We introduce MambaVision, which is a novel hybrid Mamba-Transformer model. The hierarchical MambaVision achieves a new SOTA Pareto front on the ImageNet-1K dataset in terms of Top-1 and image throughput.
Figure 2: The architecture of hierarchical MambaVision models. The first two stages use residual convolutional blocks for fast feature extraction. Stages 3 and 4 employ both MambaVision and Transformer blocks. Specifically, given $N$ layers, we use $\frac{N}{2}$ MambaVision and MLP blocks, which are followed by additional $\frac{N}{2}$ Transformer and MLP blocks. The Transformer blocks in the final layers allow for recovering lost global context and capturing long-range spatial dependencies.
@@ 2 Related work
Conv-Based. CNNs have been the cornerstone of computer vision since the introduction of AlexNet [16]. Recent efforts have focused on modernizing CNN architectures with Transformer-inspired principles. ConvNeXt [23] demonstrated competitive performance with Transformers by redesigning ResNet [12] with increased width, larger kernels, and layer normalization. RegNetY [26] introduced systematic network design through design space analysis, while EfficientNetV2 [27] leveraged neural architecture search and progressive learning for better efficiency trade-offs. Despite strong performance, these CNN models inherently lack the global receptive field needed for capturing long-range dependencies.
Transformer-Based. ViTs [5] marked a significant shift in computer vision by introducing self-attention mechanisms for enlarged receptive fields. However, ViTs initially lacked the inherent advantages of CNNs and required extensive training data. To address these limitations, DeiT [28] introduced distillation-based training for improved accuracy on smaller datasets, while Swin Transformer [21] proposed a hierarchical architecture using shifted windows for self-attention, effectively balancing local and global context. Models such as Twins [3] and PVT [32] further enhanced efficiency through spatially separable self-attention and hierarchical structures with patch embedding. Despite these advances, the quadratic complexity of self-attention operations in these models continued to pose efficiency challenges.
Conv-Transformer. The complementary strengths of CNNs and ViTs inspired hybrid architectures. CoAT [36] and CrossViT [1] demonstrated enhanced feature learning by combining convolutions with self-attention, while NextViT [17] systematically incorporated CNN-like processing into Transformers. Recent efforts like EfficientFormer [18] and FasterViT [10] focused on optimizing efficiency-accuracy trade-offs, achieving competitive performance with high throughput through carefully designed hybrid architectures.
Mamba-Based. Since the introduction of Mamba, a number of efforts have been proposed to leverage its capability for vision applications. Vim [40] introduced a bidirectional SSM formulation that processes tokens in both forward and backward directions to capture global context and improve spatial understanding. However, this bidirectional approach faces significant limitations: it increases computational overhead, slows down training and inference times, and struggles to effectively combine information from multiple directions without losing global context. In contrast, MambaVision achieves superior results using a single forward pass with our redesigned Mamba block, demonstrating significantly better ImageNet Top-1 accuracy and throughput. VMamba [20] proposed a generic Mamba-based vision backbone featuring a Cross-Scan Module (CSM). This module implements a four-way selective scan methodology (from upper-left and lower-right to opposite directions) to integrate information from surrounding tokens and capture global context. VMamba also incorporates architectural modifications like depth-wise convolutions and a hierarchical multi-resolution structure. While the CSM module is designed for vision tasks, its receptive field remains constrained by the cross-scan paths. MambaVision offers several advantages over VMamba: our mixer design is simpler yet captures both short and long-range dependencies more effectively, we employ CNN-based layers for fast feature extraction rather than using uniform block structures across all stages, and we achieve superior performance with significantly higher throughput. EfficientVMamba [25] uses SSMs for larger resolutions and CNNs for lower ones, while MambaVision takes the opposite approach with CNNs at higher resolutions and SSM/self-attention at lower ones, leading to significantly better accuracy and throughput. Similarly, while SiMBA [24] addresses Mamba’s stability through EinFFT channel modeling, it doesn’t fully address spatial understanding limitations, resulting in lower performance compared to MambaVision’s comprehensive design.
3 Methodology
@@ 3.1 Macro Architecture
In this section, we introduce MambaVision which is our proposed novel architecture with SOTA performance on ImageNet-1K dataset. As illustrated in Fig. 2, MambaVision has a hierarchical architecture consisting of $4$ different stages. The first two stages consist of CNN-based layers for fast feature extraction at higher input resolutions, while stage $3$ and $4$ include the proposed MambaVision and Transformer blocks. Specifically, given an image of size $H\times W\times 3$, the input is first converted into overlapping patches with size $\frac{H}{4}\times\frac{W}{4}\times C$ and projected into a $C$ dimensional embedding space by the stem which consists of two consecutive $3\times 3$ CNN layers with stride of $2$. The downsampler in between stages consists of a $3\times 3$ CNN layer with stride $2$ which reduces the image resolution by half. Furthermore, the CNN blocks in stages $1$ and $2$ follow a generic residual block formulation according to the following
@@
GELU and BN denote Gaussian Error Linear Unit activation function [14] and batch normalization [15], respectively. Please see the supplementary materials for further details regarding MambaVision macro architecture.
@@
dt = torch.exp(torch.rand(self.dim//2) * (math.log(dt_max) - math.log(dt_min)) + math.log(dt_min))
@@
3.2 Micro Architecture
3.2.1 Mamba Preliminaries
In Mamba, a 1D continuous input $x(t)\in\mathbb{R}$ is transformed into $y(t)\in\mathbb{R}$ via a learnable hidden state $h(t)\in\mathbb{R}^{M}$ with parameters $\bm{A}\in\mathbb{R}^{M\times M}$, $\bm{B}\in\mathbb{R}^{M\times 1}$ and $\bm{C}\in\mathbb{R}^{1\times M}$ according to
Discretization
The continuous parameters $\bm{A}$, $\bm{B}$ and $\bm{C}$ in the above formulation are further converted into discrete parameters for better computational efficiency [8]. Specifically, assuming a timescale $\Delta$, a zero-order hold rule can be applied to obtain discrete parameters $\bm{\bar{A}}\in\mathbb{R}^{M\times M}$, $\bm{\bar{B}}\in\mathbb{R}^{M\times 1}$ and $\bm{\bar{C}}\in\mathbb{R}^{1\times M}$ according to
$\bar{\bm{A}}=\exp{(\Delta\bm{A})},$ (3)
$\bar{\bm{C}}=\bm{C},$
The Eq. 2 can then be expressed with discrete parameters as
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$ (4)
$y(t)=\bar{\bm{C}}h(t),$
$\begin{split}\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{T-1}\bm{\overline{B}}),\\ \quad\bm{y}&=\bm{x}*\bm{\overline{K}},\\ \end{split}$ (5)
Selectivity
3.2.2 Layer Architecture
$\begin{array}[]{l}\hat{{X}}^{n}=\text{Mixer}(\text{Norm}({X}^{n-1}))+{X}^{n-1},\\ {X}^{n}=\text{MLP}(\text{Norm}(\hat{{X}}^{n}))+\hat{{X}}^{n},\\ \end{array}$ (6)
@@ 3.2.2 Layer Architecture
As shown in Fig. 3, we redesigned the original Mamba mixer to make it more suitable for vision tasks. First, we propose to replace the causal convolution with regular convolution, since it limits the influence to one direction, which is unnecessary and restrictive for vision tasks. In addition, we added a symmetric branch without SSM, consisting of an additional convolution and Sigmoid Linear Unit (SiLU) [6] activation, to compensate for any content lost due to the sequential constraints of SSMs. We then concatenate the output of both branches and project it via a final linear layer. This combination ensures that the final feature representation incorporates both the sequential and spatial information, leveraging the strengths of both branches. We note that the output of each branch is projected into an embedding space with size $\frac{C}{2}$ (i.e. half the size of original embedding dimension) to maintain similar number of parameters to the original block design. Given an input $X_{in}$, the output of MambaVision mixer $X_{out}$ is computed according to
$\begin{array}[]{l}{X_{1}}=\text{Scan}(\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}})))),\\ {X_{2}}=\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}}))),\\ {X_{out}}=\text{Linear}(\frac{C}{2},C)(\text{Concat}(X_{1},X_{2})),\end{array}$ (7)
$\text{Linear}(C_{in},C_{out})(\cdot)$ denotes a linear layer with $C_{in}$ and $C_{out}$ as input and output embedding dimensions, Scan is the selective scan operation as in [7] and $\sigma$ is the activation function for which SiLU is used. In addition, Conv and Concat represent 1D convolution and concatenation operations. In Algorithm 1, we present a PyTorch-like pseudo-code for MambaVision mixer. In general, our proposed modification leads to richer feature representations, better generalization, and improved performance on computer vision tasks. We have also experimentally validated the effectiveness of each of our design choices in Sec. 5.3.
@@
$Q,K,V$ denote query, key and value respectively and $d_{h}$ is the number of attention heads. In addition, our framework allows for computing the attention in a windowed manner similar to previous efforts [21, 22] (see the supplementary materials for window size ablation study).
4 Experiments
Image classification experiments are conducted on the ImageNet-1K dataset [4]. We followed the standard training recipe of previous efforts [21, 37, 11] to allow for a comparable analysis of performance across different models. Specifically, all models have been trained for 300 epochs using 32 A100 GPUs. The self-attention formulation in stages 3 and 4 of all MambaVision variants use a window size of 14 and 7, respectively. For detailed training configurations, please see the provided anonymous code repository. To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection, instance segmentation, and semantic segmentation tasks using the MS COCO dataset [19] and ADE20K dataset [39], respectively. Specifically, for object detection and instance segmentation, we used Cascade Mask-RCNN [13] head with hyperparameters such as ×3 LR schedule. For semantic segmentation, we used a UperNet network [34] head and 8 A100 GPUs for all experiments.
Table 1: Comparison of classification benchmarks on ImageNet-1K dataset [4]. Image throughput is measured on A100 GPU with a batch size of 128.
Model Image Size #Params FLOPs Throughput Top-1